# Embedding ample semigroups as $(2,1,1)$-subalgebras of inverse semigroups

### v2

@@ 1. Motivation and preliminaries
where $\rho_{s}:Tss^{-1}\longrightarrow Ts^{-1}s$ is the partial bijection defined in Theorem 1.1. Denoting $\rho_{s}|_{Dom\rho_{s}\cap Y}$, equivalently $\rho_{s}|_{Tss^{-1}\cap Y}$, by $\sigma_{s}$, we introduce a map:
$\hat{\rho}_{{}_{Y}}:S\longrightarrow\mathcal{I}_{Y}\text{ by }(s)\hat{\rho}_{{}_{Y}}=\sigma_{s}.$
Clearly, $\hat{\rho}_{{}_{Y}}$ is well-defined. Also, it is known to be a homomorphism when $S$ is finite; see [1]. The following lemma shows that $\hat{\rho}_{{}_{Y}}$ is a homomorphism even if $S$ is not finite.
@@
With the notations defined above, $\hat{\rho}_{{}_{Y}}$ is a homomorphism.
@@
We need to show that $\sigma_{s}\sigma_{t}=\sigma_{st}$, for all $s,t\in S$. By Theorem 1.1 we have $\rho_{st}=\rho_{s}\rho_{t}$, also requiring that
@@
$\sigma_{s}\sigma_{t}=\left(\rho_{s}|_{Tss^{-1}\cap Y}\right)\left(\rho_{t}|_{Ttt^{-1}\cap Y}\right).$
whence,
$(x)\sigma_{st}=(x)(\rho_{s}\rho_{t})=(x)(\sigma_{s}\sigma_{t}),\,\forall\,x\in Dom(\sigma_{s}\sigma_{t})\cap Dom(\sigma_{st}).$
So, the proof will be accomplished if we show that $Dom(\sigma_{s}\sigma_{t})=Dom(\sigma_{st})$, that is,
$(Tss^{-1}\cap Y)\rho_{s}\cap(Ttt^{-1}\cap Y)]\rho^{-1}_{s}=(Ts^{-1}s\cap Ttt^{-1})\rho^{-1}_{s}\cap Y.$
To this end, observe that
@@
where the ‘if’ part of the last implication follows by observing that
@@
This implies that $Dom(\sigma_{s}\sigma_{t})=Dom(\sigma_{st})$, and the proof is complete. ∎
In particular, the subsemigroup $S$ is a right $S$-invariant subset of $T$, giving a homomorphism $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$, as defined above.
@@
If $S$ is left ample in $T$, then $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$ is a monomorphism. Moreover, $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$.
@@
To prove the first part of the proposition, we only need to show that $\hat{\rho}_{{}_{S}}$ is injective. Assume, for this purpose, that
$\sigma_{s}=(s)\hat{\rho}_{{}_{S}}=(t)\hat{\rho}_{{}_{S}}=\sigma_{t},\;\text{i.e. }\,\rho_{s}|_{Tss^{-1}\cap S}=\rho_{t}|_{Ttt^{-1}\cap S},$
for some $s,t\in S$. Then, $ss^{-1}\in Tss^{-1}\cap S$ and $tt^{-1}\in Ttt^{-1}\cap S$, because $S$ is left ample, and we may calculate from $Tss^{-1}\cap S=Ttt^{-1}\cap S$ (viz. the equality of domains of $\sigma_{s}$ and $\sigma_{t}$) that
@@
Consequently, we have $s\leq t\leq s$ in $T$, yielding $s=t$. Hence, $\hat{\rho}_{{}_{S}}$ is an injection.
@@
$Sxx^{-1}\subseteq Txx^{-1}\cap S,\ \forall\,x\in S.$
Then observe that, for all $y\in T,\,x\in S$, we have
$y(xx^{-1})\in Txx^{-1}\cap S\implies y(xx^{-1})=y(xx^{-1})(xx^{-1})\in Sxx^{-1}.$
Thus, for every $x\in S$, one has
$Sxx^{-1}=Txx^{-1}\cap S.$ (2.1)
This implies that the domain and codomain of $\sigma_{s}$ are indeed
$Sss^{-1}\text{ and }Sss^{-1}s=Ss\,\,(\subseteq Ss^{-1}s\cap S),$ (2.2)
respectively. Let us next note that
$(\rho_{{}_{T}})^{-1}\circ\hat{\rho}_{{}_{S}}:(S)\rho_{{}_{T}}\longrightarrow(S)\hat{\rho}_{{}_{S}}$
defined by
$\rho_{s}\longmapsto\rho_{s}|_{Dom\rho_{s}\cap S}=\sigma_{s},\,s\in S,$
is an isomorphism. Now, because $\sigma_{x}\sigma_{x}^{-1}$ is the identity on $Sxx^{-1}=Txx^{-1}\cap S$ and $\rho_{xx^{-1}}$ is the identity on $Txx^{-1}$ ($Dom\rho_{x}$), where $x^{-1}$ is the inverse of $x$ in $T$, we have
$\sigma_{x}\sigma_{x}^{-1}=\rho_{xx^{-1}}|_{Dom\rho_{xx^{-1}}\cap S}.$
Finally, since $xx^{-1}\in S$, it follows from the above equality that
$\sigma_{x}\sigma_{x}^{-1}=(\rho_{xx^{-1}})\left((\rho_{{}_{T}})^{-1}\circ\hat{\rho}_{{}_{S}}\right)=\left((\rho_{xx^{-1}})(\rho_{{}_{T}})^{-1}\right)\hat{\rho}_{{}_{S}}\in(S)\hat{\rho}_{{}_{S}}.$
Hence $\sigma_{x}\sigma_{x}^{-1}\in(S)\hat{\rho}_{{}_{S}}$, implying that $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$. ∎
@@
Let $S$ be left ample in $T$. Also, let $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$ and $\sigma_{s}:Sss^{-1}\longrightarrow Ss$ be as considered in Proposition 2.2. Then, considering the isomorphism
$\theta=(\rho_{{}_{T}}|_{S})^{{}^{-1}}\circ\hat{\rho}_{{}_{S}}:(S)\rho_{{}_{T}}\longrightarrow\mathcal{I}_{S},\ \text{ defined by }\ \rho_{s}\longmapsto\sigma_{s},\forall\,s\in S,$
we have, for all $s\in S$,
@@
Using (2.1), let us first calculate
@@
Then, we have
$(\rho_{s}\rho_{s}^{-1})\theta=(\rho_{s}\rho_{s^{-1}})\theta=(\rho_{ss^{-1}})\theta=\left((\rho_{ss^{-1}})(\rho_{{}_{T}}|_{S})^{{}^{-1}}\right)\hat{\rho}_{{}_{S}}=(ss^{-1})\hat{\rho}_{{}_{S}}=\sigma_{ss^{-1}}=\sigma_{s}\sigma_{s}^{-1},$
completing the proof. ∎
For any left ample semigroup $S$, it follows from the above corollary that for all $x\in S$,
$\left(\sigma_{x}\sigma_{x}^{-1}\right)\hat{\rho}_{{}_{S}}^{-1}=\left(\rho_{x}\rho_{x}^{-1}\right){\rho}_{{}_{T}}^{-1}=xx^{-1}\in S.$
The following theorem is essentially a restatement of Proposition 2.2.
@@
Let $U$ be a left ample semigroup (with some associated inverse semigroup $T$). Then $\hat{\rho}_{{}_{U}}:U\longrightarrow\mathcal{I}_{U}$, given by $x\mapsto\sigma^{\prime}_{x}$, where $\sigma^{\prime}_{x}:Uxx^{-1}\longrightarrow Ux$ is defined by $(z)\sigma^{\prime}_{x}=zx$ for all $z\in Uxx^{-1}$, is a monomorphism. Moreover, $(U)\hat{\rho}_{{}_{U}}$ is left ample in $\mathcal{I}_{U}$.
@@
The morphisms maps used in the proof are shown in Figure 1 (the morphism $\theta$ was defined in Corollary 2.3). Let $T$ be an (associated) inverse semigroup admitting a monomorphism
$\phi:U\longrightarrow T,$
such that $(U)\phi=S$ is left ample in $T$. Let $\phi^{\prime}:U\longrightarrow S$ be the isomorphism defined by $x\mapsto(x)\phi$, and $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$, defined by
$(x)\phi^{\prime}\longmapsto\sigma_{(x)\phi^{\prime}},\,\forall\,x\in U,$
be the embedding given by Proposition 2.2. Also, let $\hat{\phi}=\phi^{\prime}\circ\hat{\rho}_{{}_{S}}:U\longrightarrow\mathcal{I}_{S}$. Then, defining $s=(x)\phi^{\prime}$, and using Corollary 2.3, we obtain
@@
where the last equality is due to the identification: $(x)\phi^{\prime}((x)\phi^{\prime})^{-1}=(xx^{-1})\phi^{\prime}$. Thus, we may write
$\{(x)\hat{\phi}[(x)\hat{\phi}]^{-1}\}(\hat{\phi})^{-1}=xx^{-1}=\{[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}\in U.\vskip 2.0pt$ (2.4)
@@
Now, for an arbitrary partial bijection $\beta:A\longrightarrow B$ in $\mathcal{I}_{S}$, consider the map
@@
$(a)(\phi^{\prime})^{-1}\longmapsto((a)\beta)(\phi^{\prime})^{-1},\,\forall\,a\in A.$ (2.5)
Clearly, $\gamma$ is an element of $\mathcal{I}_{U}$, and it is an easy exercise to prove that $\chi:\mathcal{I}_{S}\longrightarrow\mathcal{I}_{U}$ given by $\beta\mapsto\gamma$ is an isomorphism. Moreover, by Remark 1.6, $((U)\hat{\phi})\chi$ is left ample in $\mathcal{I}_{U}$, proving the second part of the proposition.
To complete the proof, let us consider the monomorphism
$\pi=\hat{\phi}\circ{\chi}\colon U\longrightarrow\mathcal{I}_{U}.$
The aim is to show that for all $x\in U$, $(x)\pi=\sigma^{\prime}_{x}$ where $\sigma^{\prime}_{x}$ is as given in the statement of the theorem. Recall that the domain and codomain of $\sigma_{(x)\phi}$ ($\sigma_{s}$) are, respectively, $(S)[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}$ and $(S)[(x)\phi^{\prime}]$. Hence, by the definitions of $\pi$ and $\chi$, we have
$(x)\pi=\left(\sigma_{(x)\phi}\right)\chi:\{S[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}\longrightarrow\{S[(x)\phi^{\prime}]\}(\phi^{\prime})^{-1},\vskip 2.0pt$
@@
$\{S[(x)\phi^{\prime}]\}(\phi^{\prime})^{-1}=Ux,$
and by (2.4), we have $\{[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}=xx^{-1}$, whence
$\{S[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}=Uxx^{-1}.$
Also, taking $z=uxx^{-1},u\in U$, we may calculate,
$\begin{split}(z)[(x)\pi]&=((uxx^{-1})[(x)\hat{\phi}\circ\chi]\\ &=\left(uxx^{-1}\right)\left(\left(\sigma_{(x)\phi^{\prime}}\right)\chi\right)\\ &=((uxx^{-1})\phi^{\prime}(x)\phi^{\prime})\phi^{\prime-1}\\ &=((uxx^{-1}x)\phi^{\prime})\phi^{\prime-1}\\ &=uxx^{-1}x\\ &=zx.\end{split}$
So, we may conclude that
$(x)\pi:Uxx^{-1}\longrightarrow Ux,$
is defined by $(z)[(x)\pi]=zx$ for all $z=uxx^{-1}\in Uxx^{-1}$, where $u\in U$. ∎
@@
Let $S$ be left as well as right ample in $T$ with $Ts^{-1}s\cap S=Ss$ for all $s\in S$. Then $(S)\hat{\rho}_{{}_{S}}$ is left as well right ample in $\mathcal{I}_{S}$.

### v3

@@ 2. Answering the embedding problem for some finite ample semigroups
where $\rho_{s}:Tss^{-1}\longrightarrow Ts^{-1}s$ is the partial bijection defined in Theorem 1.1. Denoting $\rho_{s}|_{Dom\rho_{s}\cap Y}$, equivalently $\rho_{s}|_{Tss^{-1}\cap Y}$, by $\sigma^{Y}_{s}$, we introduce a map:
$\sigma^{Y}:S\longrightarrow\mathcal{I}_{Y}\text{ by }(s)\sigma^{{}_{Y}}=\sigma^{Y}_{s}.$ (2.1)
Clearly, $\sigma^{Y}$ is well-defined. Also, it is known to be a homomorphism when $S$ is finite; see [3]. The following lemma shows that the finiteness condition is, nevertheless, not necessary.
@@
Let $Y$ be a right $S$-invarient subset of $T$. Then, with the notations defined above, $\sigma^{Y}$ is a homomorphism.
@@
We need to show that $\sigma^{Y}_{s}\sigma^{Y}_{t}=\sigma^{Y}_{st}$, for all $s,t\in S$. By Theorem 1.1 we have $\rho_{st}=\rho_{s}\rho_{t}$, whence
@@
$\sigma^{Y}_{s}\sigma^{Y}_{t}=\left(\rho_{s}|_{Tss^{-1}\cap Y}\right)\left(\rho_{t}|_{Ttt^{-1}\cap Y}\right).$
Particularly, we note that
$\begin{split}Dom(\sigma^{Y}_{st})&=(Ts^{-1}s\cap Ttt^{-1})\rho^{-1}_{s}\cap Y\\ Dom(\sigma^{Y}_{s}\sigma^{Y}_{t})&=[(Tss^{-1}\cap Y)\rho_{s}\cap(Ttt^{-1}\cap Y)]\rho_{s}^{-1}.\end{split}$
Now, using the fact that $\sigma^{Y}_{x}$ is a restrictions of $\rho_{x}$, for all $x\in S$, we may write for all $x\in Dom(\sigma^{Y}_{st})\cap Dom(\sigma^{Y}_{s}\sigma^{Y}_{t})$:
$(x)\sigma^{Y}_{st}=(x)\rho_{st}=(x)(\rho_{s}\rho_{t})=(x)(\sigma^{Y}_{s}\sigma^{Y}_{t}).$
Thus, the proof will be accomplished if we show that $Dom(\sigma^{Y}_{s}\sigma^{Y}_{t})=Dom(\sigma^{Y}_{st})$. To this end, observe that
@@
[…]
@@
Hence $Dom(\sigma^{Y}_{s}\sigma^{Y}_{t})=Dom(\sigma^{Y}_{st})$, and the proof is complete. ∎
@@
Let $S$ be a subsemigroup of an inverse semigroup $T$. Then the map $\sigma^{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$, given by (2.1), is a homomorphism.
@@
Follows from the fact that every subsemigroup $S$ of an inverse semigroup $T$ is a right $S$-invariant subset of $T$. ∎
In what follows, dropping the superscript, we shall simply write $\sigma$ instead of $\sigma^{{}_{S}}$. Although the first part of the following proposition also follows from [5] Theorem 6.2 (iv), our essentially direct scheme of proof will be instructive in the sequel.
Proposition 2.3.
If $S$ is left ample in $T$, then $\sigma:S\longrightarrow\mathcal{I}_{S}$, as given by Proposition 2.2, is a monomorphism. Also, for every $s\in S$ the domain and image of $(s)\sigma=\sigma_{s}$ are the sets $Sss^{-1}$ and $Ss$, respectively. Moreover, $(S)\sigma$ is left ample in $\mathcal{I}_{S}$.
Proof.
To prove the first part of the proposition, we only need to show that $\sigma$ is injective. Assume, for this purpose, that
$\sigma_{s}=(s)\sigma=(t)\sigma=\sigma_{t},\;\text{i.e. }\,\rho_{s}|_{Tss^{-1}\cap S}=\rho_{t}|_{Ttt^{-1}\cap S},$
for some $s,t\in S$. Because $S$ is left ample, we have $ss^{-1}\in Tss^{-1}\cap S$ and $tt^{-1}\in Ttt^{-1}\cap S$. Consequently, we may claim, using $Tss^{-1}\cap S=Ttt^{-1}\cap S$ (viz. the equality of domains of $\sigma_{s}$ and $\sigma_{t}$), that
@@
Now, it follows from the definition of order on $T$ that $s\leq t\leq s$. Thus we get $s=t$, whence $\sigma$ is injective.
@@
$Sss^{-1}\subseteq Tss^{-1}\cap S,\ \forall\,s\in S.$
Then observe that, for all $x\in T,\,s\in S$,
$x(ss^{-1})\in Tss^{-1}\cap S\implies x(ss^{-1})=x(ss^{-1})(ss^{-1})\in Sss^{-1}.$
Hence, for every $s\in S$ one has
$Sss^{-1}=Tss^{-1}\cap S.$ (2.2)
This implies that for every $s\in S$, the domain and image of $\sigma_{s}$ are, respectively, the sets
$Sss^{-1}\text{ and }Sss^{-1}s=Ss\,\,(\subseteq Ss^{-1}s\cap S).$
To prove the last part of the proposition, we first consider the isomorphism $\rho^{\prime}:T\rho\longrightarrow T$ defined by $\rho_{x}\longmapsto x$ for all $x\in T$. Then, one may define an isomorphism
$\rho^{\prime}|_{S\rho}\circ\sigma:S\rho\longrightarrow S\sigma$
given by
$\rho_{s}\longmapsto\sigma_{s},\,\forall\,s\in S.$
Next, because $\sigma_{s}\sigma_{s}^{-1}$ is the identity on $Sss^{-1}=Tss^{-1}\cap S$ and $\rho_{ss^{-1}}$ is the identity on $Tss^{-1}$, we have
$\sigma_{s}\sigma_{s}^{-1}=\rho_{ss^{-1}}|_{Dom\rho_{ss^{-1}}\cap S}.$
Now, since $ss^{-1}\in Dom\rho_{ss^{-1}}\cap S$, it follows from the above equality that
$\sigma_{s}\sigma_{s}^{-1}=(\rho_{ss^{-1}})\left(\rho^{\prime}\circ\sigma\right)=\left((\rho_{ss^{-1}})\rho^{\prime}\right)\sigma\in S\sigma.$
implying that $S\sigma$ is left ample in $\mathcal{I}_{S}$. ∎
@@
Let $S$ be left ample in $T$. Also, let $\sigma_{s}=(s)\sigma$ and $\rho_{s}=(s)\rho$ be as considered in Proposition 2.3. Then, considering the isomorphism
$\theta:(S)\rho\longrightarrow\mathcal{I}_{S},\text{ defined by }\ \rho_{s}\longmapsto\sigma_{s},\forall\,s\in S$
from Proposition 2.3, we have, for all $s\in S$,
@@
Clearly, we have
$(\rho_{s}\rho_{s}^{-1})\theta=(\rho_{s}\rho_{s^{-1}})\theta=(\rho_{ss^{-1}})\theta=\sigma_{ss^{-1}}.$
Also, using (2.2), we can calculate:
@@
This completes the proof. ∎
With $S$ replaced with $S_{1}$, Figure 1 shows all maps used in Proposition 2.3, Corollary 2.4.
Remark 2.5.
For any left ample semigroup $S$, it follows from the above corollary that
$\left(\sigma_{s}\sigma_{s}^{-1}\right)\sigma^{\prime}=ss^{-1}=\left(\rho_{s}\rho_{s}^{-1}\right)\rho^{\prime},\,\forall\,s\in S,$
with $\rho^{\prime}$ introduced in the proof of Proposition 2.3 and $\sigma^{\prime}$ being defined analogously.
In Proposition 2.3 and its corollary, $S$ was left ample in $T$. In the following theorem $S$ is not necessarily contained in its associated inverse semigroup. (In fact, this result relates more closely to Theorem 6.2 (iv) of [5]; cf. the paragraph before Proposition 2.3.)
@@
Let $S$ be a left ample semigroup. Then, for every $x\in S$ the mapping $\hat{\rho}_{x}:Sxx^{-1}\longrightarrow Sx$ defined by $(z)\hat{\rho}_{x}=zx$, for all $z\in Sxx^{-1}$, is a bijection. Furthermore, the function $\hat{\rho}:S\longrightarrow\mathcal{I}_{S}$ given by $x\longmapsto\hat{\rho}_{x}$ is a monomorphism, and $(S)\hat{\rho}$ is left ample in $\mathcal{I}_{S}$.
@@
To help the reader, the morphisms used in the proof are shown in Figure 1. Let $T$ be an (associated) inverse semigroup admitting a monomorphism
$\phi:S\longrightarrow T,$
such that $(S)\phi=S_{1}$ is left ample in $T$. Let $\phi^{\prime}:S\longrightarrow S_{1}$ be the isomorphism defined by $x\longmapsto(x)\phi$, and $\sigma:S_{1}\longrightarrow\mathcal{I}_{S_{1}}$, defined by
$(x)\phi^{\prime}\longmapsto\sigma_{(x)\phi^{\prime}},\,\forall\,x\in S,$
be the monomorphism given by Proposition 2.3. Particularly, $S_{1}\sigma$ is left ample in $\mathcal{I}_{S_{1}}$.
Let $\hat{\phi}=\phi^{\prime}\circ\sigma:S\longrightarrow\mathcal{I}_{S_{1}}$, and $x$ be an arbitrary element of $S$. Then, defining $s=(x)\phi^{\prime}$, and using Corollary 2.4, we may calculate,
@@
Thus, we may write
$(x\hat{\phi})(x\hat{\phi})^{-1}]\hat{\phi}^{-1}=[(x\phi^{\prime})(x\phi^{\prime})^{-1}]\phi^{\prime-1}=xx^{-1}\in S.\vskip 2.0pt$ (2.3)
@@
Now, for an arbitrary partial bijection $\beta:A\longrightarrow B$ in $\mathcal{I}_{S_{1}}$, consider the map
@@
$(a)(\phi^{\prime})^{-1}\longmapsto(a\beta)(\phi^{\prime})^{-1},\,\forall\,a\in A.$ (2.4)
Clearly, $\gamma$ is an element of $\mathcal{I}_{S}$, and it is a routine to verify that $\chi:\mathcal{I}_{S_{1}}\longrightarrow\mathcal{I}_{S}$ given by $\beta\longmapsto\gamma$ is an isomorphism. Moreover, by Remark 1.6, $(S\hat{\phi})\chi=(S_{1}\sigma)\chi$ is left ample in $\mathcal{I}_{S}$.
To complete the proof, let us consider the monomorphism $\pi=\hat{\phi}\circ{\chi}\colon S\longrightarrow\mathcal{I}_{S}$. The aim is to show that $\pi=\hat{\rho}$, i.e. for all $x\in S$, $(x)\pi=\hat{\rho}_{x}$ where $\hat{\rho}_{x}$ is as given in the statement of the theorem. Recall from Proposition 2.3 that the domain and image of $\sigma_{(x)\phi}$ ($\sigma_{s}$) are, respectively, $(S_{1})(x\phi^{\prime})(x\phi^{\prime})^{-1}$ and $(S_{1})(x\phi^{\prime})$. Hence, by the definitions of $\pi$ and $\chi$, we have
$(x)\pi=\left(\sigma_{(x)\phi}\right)\chi:\{S_{1}(x\phi^{\prime})(x\phi^{\prime})^{-1}\}(\phi^{\prime})^{-1}\longrightarrow\{S_{1}(x\phi^{\prime})\}(\phi^{\prime})^{-1},\vskip 2.0pt$
@@
$\{S_{1}(x\phi^{\prime})\}(\phi^{\prime})^{-1}=Sx,$
and by (2.3), we have
$\{S_{1}(x\phi^{\prime})(x\phi^{\prime})^{-1}\}(\phi^{\prime})^{-1}=Sxx^{-1}.$
Also, taking $z=uxx^{-1},u\in S$, we may calculate, using (2.4),
[…]
$\begin{split}&=((uxx^{-1}x)\phi^{\prime})\phi^{\prime-1}\\ &=uxx^{-1}x\\ &=zx.\end{split}$
So, we conclude that
$(x)\pi:Sxx^{-1}\longrightarrow Sx,$
is defined by $(z)(x\pi)=zx$ for all $z=uxx^{-1}\in Sxx^{-1}$ ($u\in S$), hence the proof. ∎
Before stating our main result, let us note that using the dual notion of left $S$-invariance, one can similarly prove that any right ample subsemigroup $S$ of an inverse semigroup $T$ has an anti-isomorphic copy in the symmetric inverse semigroup $\mathcal{I}_{S}$ given by the anti-monomorphism,
$\delta:S\longrightarrow\mathcal{I}_{S},\hskip 5.0pt\text{given by }s\longmapsto{\delta_{s}}={\lambda_{s}}|_{s^{-1}sT\,\cap\,S},\,\forall\,s\in S,$
where $\lambda_{s}\in\mathcal{I}_{T}$ is the partial bijection defined in Theorem 1.2. Moreover, $s^{-1}sT\cap S=s^{-1}sS$, and we have the following dual of Theorem 2.6.
@@
For any right ample semigroup $S$, there exists an anti-monomorphism,
[…]
where the partial bijection,
$\hat{\lambda}_{x}:x^{-1}xS\longrightarrow xS$
is given by $(z)\hat{\lambda}_{x}=xz$, for all $z\in x^{-1}xS$. Moreover, $(S)\hat{\lambda}$ is left ample in $\mathcal{I}_{S}$.
We now prove the main theorem of this section.
Figure 2.
Theorem 2.8.
Let $S$ be a finite ample semigroup such that for all $e\in E(S)$ there exists a bijection between the principal ideals $eS$ and $Se$. Then $S$ is ample in, and hence a $(2,1,1)$-subalgebra of, $\mathcal{I}_{S}$.